A normal distribution has mean 50 and standard deviation 4. Approximately what proportion

of this normal distribution lies above 58?
Answer #1

P[X>58]

58 - 50 = P(Z >

=P[Z>2]

= 1-P2<2

=1-0.9772...........................by using Z table

=0.0228